OpenStudy (anonymous):

write the following quandratic function in vertex form. y=x^2+2x-2

OpenStudy (anonymous):

is this the same mathdummy4 that was asking about the closed form of \(\int e^{x^2}dx\)?

OpenStudy (anonymous):

lol ya...kinda a long story..

OpenStudy (anonymous):

to find the first coordinate of the vertex , use \(-\frac{b}{2a}\) in this case you get \(-\frac{2}{2\times 1}=-1\)

OpenStudy (anonymous):

second coordinate of the vertex is what you get when you replace \(x\) by \(-1\) namely \[(-1)^2+2\times (-1)-2=1-2-2=-3\]

OpenStudy (anonymous):

vertex form is therefore \(y=(x+1)^2-3\) not sure what the closed form is though...

OpenStudy (anonymous):

that all im looking for..thanks satellite! you sure are a life-saver